Table: Production Possibilities for Tractors and Crude Oil

Country Tractors Crude Oil
(thousands of barrels)
United States 80 40
Mexico 60 180

Reference: Ref 8-5 Table: The Production Possibilities for Tractors and Crude Oil


(Table: The Production Possibilities for Tractors and Crude Oil) Use Table: The Production Possibilities for Tractors and Crude Oil. The opportunity cost of _____ is _____ in the United States as/than in Mexico.

Select one:

a. tractors; the same

b. tractors; lower

c. crude oil; lower

d. tractors; higher
Answer #1

USA opportunity cost: For tractor : 40/80 =0.5

For crude oil: 80/40=2

Mexico opportunity cost: For tractor : 180/60= 3

For crude oil: 60/180=0.33

hence USA to focus on tractor and Mexico on crude oil.

Hence correct answer is option: b.

The opportunity cost of tractor is lower in USA is less than in Mexico.